OpenStudy (anonymous):

d/dx integral (0,x^5) sin^2t(dt)

OpenStudy (anonymous):

To make sure I understand you correctly, is this your question? \[ \frac{d}{dx}\int_0^{x^5}sin^2t\ dt \]

OpenStudy (anonymous):

0 and x^5 spots are switched

OpenStudy (anonymous):

Okay, so \[\frac{d}{dx}\int_{x^5}^0sin^2t\ dt\] Just apply the chain rule. Remember, \((f\circ g)'(x)=f'(g(x))\cdot g'(x)\). In this problem you'll have \(f(x)=\int_x^0 \sin^2t\ dt\), and \(g(x)=x^5\)

OpenStudy (anonymous):

(-cos^2dt)(x^5)*5x^4???

OpenStudy (anonymous):

Look @nbousca 's instruction closely!

OpenStudy (anonymous):

f(x) = integral .... -> f '(x) = derivative of integral

OpenStudy (anonymous):

First we derive \(f\), which gives us \(f'(t)=-\sin^2t\) by the FTC. Then plug in \(g(x)\) for \(t\), then multiply by \(g'(x)\).